# On the Burst-Covering Radius of Binary Cyclic Codes

### v1

@@ VI Burst-Covering Algorithm for Binary Cyclic Codes
The time complexity of Algorithm 1 is $\mathcal{O}(r^{3}+r\cdot n)$: the term $r^{3}$ corresponds to solving the linear system in line 5; the second is the maximum number of iterations of the while loop (which is $\operatorname{ord}(g)\leqslant n$) times $r$, the degree of $g$, which determines the time required for the polynomial addition. In the case of $\operatorname{BCH}(e,m)$ codes, the time complexity becomes $\mathcal{O}(e^{3}\log^{3}n+e\cdot n\log n)$.
VII Conclusion
@@
In this paper we initiated the study over burst covering codes. We focused in particular on the rich family of cyclic codes. We showed how the burst-covering radius of cyclic codes depends on the length of runs of zeros in related LFSR sequences. In some cases we can give the exact burst-covering radius. In the specific case of BCH codes, we developed a new bound on the existence of patterns in LFSR sequences, complementing those already found in the literature. For binary BCH codes and Melas codes it appears that the upper bound on the burst-covering radius is close to the actual radius. We concluded with an efficient algorithm for burst-covering cyclic codes.
@@
Many open question remain. The immediate problem arising from the case BCH and Melas codes is the gap between the lower bound (Theorem 38) and the upper bound (Theorem 35 and Theorem 36) on the burst-covering radius. The numerical results of Table I suggest the lower bound should be significantly improved. However, that requires results on the nonexistence of patterns in LFSR sequences. To the best of our knowledge, these results are not known yet.
@@
Lemma 39.
For all integers $a,b\geqslant 1$,
$1+\frac{(2^{a}-1)(2^{b}-1)}{(2^{\gcd(a,b)}-1)2^{(a+b)/2}}>2^{(a+b)/2-\gcd(a,b)}.$
Proof:
Let $c=\gcd(a,b)$. The inequality holds if and only if
$\frac{(2^{c}-1)2^{(a+b)/2}+(2^{a}-1)(2^{b}-1)}{(2^{c}-1)2^{(a+b)/2}}>\frac{2^{(a+b)/2}}{2^{c}}\Longleftrightarrow$
$2^{c}(2^{c}-1)2^{(a+b)/2}+2^{c}2^{a+b}-2^{c}2^{a}-2^{c}2^{b}+2^{c}>2^{a+b}(2^{c}-1)\Longleftrightarrow$
$2^{c}(2^{c}-1)2^{(a+b)/2}+2^{a+b}+2^{c}>2^{a+c}+2^{b+c}$ (12)
We analyze different scenarios:
  • If $a=1$: then $c=\gcd(a,b)=1$, and the inequality becomes
    $2\cdot 2^{(b+1)/2}+2^{b+1}+2>4+2^{b+1}\Longleftrightarrow$
    $2^{(b+3)/2}>2\Longleftrightarrow(b+3)/2>1\Longleftrightarrow b>-1,$
    so the inequality (12) holds in this case.
  • If $a>1$ and $c=\gcd(a,b)\leqslant a/2$ then
    $2^{a+c}+2^{b+c}\leqslant 2^{a+a/2}+2^{b+a/2}\leqslant 2\cdot 2^{b+a/2}$
    If $a\geqslant 2$, then $a/2+1\leqslant a$, so this means $2^{a+c}+2^{b+c}\leqslant 2^{b+a/2+1}\leqslant 2^{a+b}$, and so (12) holds.
  • If $a>1$ and $c=\gcd(a,b)=a$, (12) becomes:
    $2^{a}(2^{a}-1)2^{(a+b)/2}+2^{a+b}+2^{a}>2^{2a}+2^{a+b}\Longleftrightarrow$
    $2^{3a/2+b/2}(2^{a}-1)>2^{a}(2^{a}-1)\Longleftrightarrow$
    $3a/2+b/2>a,$
    which is true for positive $a,b$.

### v2

@@ VI Burst-Covering Algorithm for Binary Cyclic Codes
The time complexity of Algorithm 1 is $\mathcal{O}(r^{3}+r\cdot n)$: the term $r^{3}$ corresponds to solving the linear system in line 5; the second is the maximum number of iterations of the while loop (which is $\leqslant\operatorname{ord}(g)\leqslant n$) times $r$, the degree of $g$, which determines the time required for the polynomial addition. In the case of $\operatorname{BCH}(e,m)$ codes, the time complexity becomes $\mathcal{O}(e^{3}\log^{3}n+e\cdot n\log n)$.
We can also give an algorithm to compute the burst-covering radius of a cyclic code with simple roots based on the characterization (7). There are $\approx 2^{r}/n$ distinct LFSR sequences with connection polynomial $g$, up to rotation. Computing the maximum run of zeros for each requires time $n$, and thus the time complexity is $\mathcal{O}(2^{r})$. In the case of $\operatorname{BCH}(e,m)$ codes, this is equal to $\mathcal{O}(n^{e})$ time.
The previous algorithm can be considerably slower than Algorithm 2 in [23], which determines the burst-error-correcting limit of a cyclic code in time $\mathcal{O}(kr^{2})$. It would be of interest to determine if the time can be improved in the covering case.
VII Connection to the critical exponent
Recall the definition of the critical exponent of a linear code:
Definition 39 ([19]).
The critical exponent of an $n,k]_{q}$ code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is the minimum number $c$ such that $\mathcal{C}$ contains $c$ codewords whose supports cover all $n$ coordinate positions.
The critical exponent of a code is related to the famously hard critical problem for matroids, posed by Crapo and Rota [11]. For binary cyclic codes, we can give a relatively straightforward upper bound for the critical exponent based on the burst-covering radius of the dual code:
Theorem 40.
Given $\mathcal{C}$ a binary [n,nr][n,n-r] cyclic code with simple roots and burst-covering radius $b$, the critical exponent of its dual code satisfies:
$c(\mathcal{C}^{\perp})\leqslant r-b+1.$
Proof:
By Corollary 24, there is a codeword $c\in\mathcal{C}^{\perp}$ such that its maximum run of zeros has length rbr-b. Then the union of the supports of $c$ and its rbr-b shifts to the right $X\cdot c\pmod{X^{n}-1},X^{2}\cdot c\pmod{X^{n}-1},\dots,X^{r-b}\cdot c\pmod{X^{n}-1}$ is $\{1,\dots,n\}$. ∎
We can instantiate Theorem 40 to the case of BCH and Melas codes to obtain the following corollary:
Corollary 41.
For $m$ and $e$ satisfying (11),
$c(\operatorname{BCH}(e,m)^{\perp})\leqslant m-1.$
For $m\geqslant 3$,
$c(\operatorname{Melas}(m)^{\perp})\leqslant m-1.$
Proof:
Simply combine Theorem 40 together with the lower bound from Theorem 38. ∎
We can compare Theorem 40 via Colloary 41 to Kung’s bound [20, Eq. (4.10)] and its improvement given in [4, Theorem 13]:
Theorem 42 (Kung’s bound [20, 4]).
If $\mathcal{C}$ is an [n,nr][n,n-r] code with minimum distance $d\geqslant 3$, then
$c(\mathcal{C}^{\perp})\leqslant r-d+3.$
Unless $\mathcal{C}$ is the Hamming code or an $[n,1]$ code with $d=n$ odd, the upper bound can be improved to rd+2r-d+2.
If we now use Theorem 42, we get
$c(\operatorname{BCH}(e,m)^{\perp})$ $\leqslant em-2e+2,$
$c(\operatorname{Melas}(m)^{\perp})$ $\leqslant 2m\phantom{-2}\qquad\text{for $m$ even,}$
$c(\operatorname{Melas}(m)^{\perp})$ $\leqslant 2m-2\qquad\text{for $m$ odd,}$
where the last two follow from the lower bound on the minimum distance of $\operatorname{Melas}(m)$ (e.g., see [28]). By comparison, the results of Corollary 41 are significantly better. If Conjecture 37 is true, it would allow us to further reduce the upper bound by approximately $m/2$.
As a final comment, it is interesting to note that while the critical exponent is invariant under permutations of the code coordinates, the upper bound of Theorem 40 depends on $b$, and thus, can change with permutations.
VIII Conclusion
@@
In this paper we initiated the study over burst covering codes. We focused in particular on the rich family of cyclic codes. We showed how the burst-covering radius of cyclic codes depends on the length of runs of zeros in related LFSR sequences. In some cases we can give the exact burst-covering radius. In the specific case of BCH codes, we developed a new bound on the existence of patterns in LFSR sequences, complementing those already found in the literature. For binary BCH codes and Melas codes it appears that the upper bound on the burst-covering radius is close to the actual radius. We then provided an efficient burst-covering algorithm for cyclic codes. Finally, we showed how the burst-covering radius of cyclic codes can help bound the critical exponent of these codes, whose computation is in general a notoriously difficult problem.
@@
Many open questions remain. The immediate problem arising from the case BCH and Melas codes is the gap between the lower bound (Theorem 38) and the upper bound (Theorem 35 and Theorem 36) on the burst-covering radius. The numerical results of Table I suggest the lower bound should be significantly improved. However, that requires results on the nonexistence of patterns in LFSR sequences. To the best of our knowledge, these results are not known yet.